# ennola_conjecture.ref.bib

### v2

@@ 1 Introduction
This is still an open problem. Ennola conjectured in [Enn91] that $\{\epsilon_{l},\epsilon_{l}-1\}$ above is a pair of fundamental units for the maximal order $\mathcal{O}$ of $\mathbb{Q}(\epsilon_{l})$. We call this the Ennola’s conjecture. He showed that his conjecture is true for $3\leq l\leq 500$ and that the unit index
@@
To show that the Ennola’s conjecture holds true is equivalent to prove $j_{l}=1$ for $\forall l\geq 3$. Louboutin obtained several results on this conjecture. For example, in [Lou17], he showed that for $l\geq 3$, the unit index $j_{l}$ is coprime to $19!$, and $j_{l}=1$ for $3\leq l\leq 5\cdot 10^{7}$. He also proved in [Lou21] that if we assume ABC conjecture is true, then the Ennola’s conjecture is true except for finitely many $l$. i.e., $j_{l}=1$ for any sufficiently large $l$. Another significant result in [Lou21] is a conditional proof of a weak form of Ennola’s conjecture that for any given integer $N\geq 2$, we have $\gcd(j_{l},N!)=1$ for $l\geq l_{N}$ effectively large enough.
@@
Theorem 1 is a statement about divisors of a unit index $j_{l}$. A similar work was done by Louboutin and Lee for another family of cubic number field. In [LL], they showed that the unit index $(\mathbb{U}_{a}:\langle-1,\epsilon_{a},\epsilon_{a}^{\prime}\rangle)$ of the group of units generated by $-1$, $\epsilon_{a}$, and $\epsilon_{a}^{\prime}$ in the group of units $\mathbb{U}_{a}$ of the ring of integers of $\mathbb{Q}(\alpha)$ is coprime to 3 for $a\geq 1$ and coprime to 6 for $1\leq a\equiv 2,3\ (\text{mod }4)$, where $\alpha$ is a root of $X^{3}-4a^{2}X+2\in\mathbb{Z}[X]$, and $\epsilon_{a},\epsilon_{a}^{\prime}$ are two distinct units in the ring of integers of $\mathbb{Q}(\alpha)$. Our result is similar to that of Louboutin and Lee, and we are concerned with a family of cubic polynomials (1) with $l\geq 3$. On the other hand, this family (1) attracts interests both of number theorists and of topologists. For topological context, see [KY].
@@ 2 Louboutin’s conjectures
One can directly compute $P_{d}(X,Y)$ for small $d\geq 1$ by finding all possible pairs of $(k,l)$, and below are some examples of $P_{d}(X,Y)$ obtained from the definition.
@@
Then for $d\in\left\{a,b,c\right\}$ and $P_{d}(X,Y)$ as in the Definition 1 below, we have
$P_{|d|}(S_{a,b}(T),S_{a,b}(1/T))=-S_{a,b}(1/T^{|d|}).$ (2)
@@
$P_{|d|}(-R_{a,b}(T),-R_{a,b}(1/T))=\begin{cases}-S_{a,b}(1/T^{|d|}),&\text{if}\ d=a\\ R_{a,b}(1/T^{|d|}),&\text{if}\ d\in\left\{b,c\right\}.\end{cases}$ (3)
@@ 3 Proof of Conjecture 1
Hence, we may regard $s_{d}(x_{1},x_{2},x_{3})$ as a polynomial of $\sigma_{1},\sigma_{2},\sigma_{3}$ for each $d\geq 4$. i.e.
$s_{d}(x_{1},x_{2},x_{3})=x_{1}^{d}+x_{2}^{d}+x_{3}^{d}=f_{d}(\sigma_{1},\sigma_{2},\sigma_{3})$ (4)
@@
$f_{d}(\sigma_{1},\sigma_{2},\sigma_{3})=\sigma_{1}f_{d-1}(\sigma_{1},\sigma_{2},\sigma_{3})-\sigma_{2}f_{d-2}(\sigma_{1},\sigma_{2},\sigma_{3})+\sigma_{3}f_{d-3}(\sigma_{1},\sigma_{2},\sigma_{3}).$ (5)
@@
$f_{d}(\sigma_{1},\sigma_{2},1)=\sigma_{1}f_{d-1}(\sigma_{1},\sigma_{2},1)-\sigma_{2}f_{d-2}(\sigma_{1},\sigma_{2},1)+f_{d-3}(\sigma_{1},\sigma_{2},1).$ (6)
@@
$-f_{d-i}(Y,X,1)=(d-i)\sum_{\begin{subarray}{c}k,l\geq 0\\ 0\leq 2k+3l\leq d-i\end{subarray}}(-1)^{k-1}\binom{k+l}{k}\binom{d-i-k-2l}{k+l}\frac{X^{k}Y^{d-i-2k-3l}}{d-i-k-2l}$ (9)
@@
$-\frac{d(d-1-2l)!}{l!\ (d-3l)!}Y^{d-3l}.$ (11)
Since we assumed $k=0$, the second term in the Case 1, which was obtained by replacing $k$ with $k-1$, cannot appear. Hence,
@@
$(-1)^{k-1}\frac{d(d-1-k)!}{k!\ (d-1-2k)!}X^{k}Y^{d-2k}$ (12)
as in Case 2 by putting $l=0$, and excluding the third term in the Case 1. That is,
@@
and the $f_{d}$ comes from Newton’s identities. Combining (4) with the discussion right after Example 2, we get the desired result. ∎
@@ 4 Proof of Conjecture 2
Table 3: $G_{a,b}(T)$ in each case.
4.1 Case 1: $a\geq 1$ odd and $b\geq 1$ odd
@@
$G_{a,b,m}(T)=F_{a,b}\left(R_{a,b}(T)+\frac{1}{mT^{m}}E_{a,b}(T),R_{-a,-b}(T)+\frac{1}{mT^{m}}E_{-a,-b}(T)\right).$ (13)
@@
$-P_{a}\left(R_{a,b}\left(\frac{1}{T}\right)-\frac{1}{mT^{m}}E_{a,b}\left(\frac{1}{T}\right),R_{a,b}(T)+\frac{1}{mT^{m}}E_{a,b}(T)\right)$ (14)
$-P_{b}\left(R_{a,b}\left(\frac{1}{T}\right)-\frac{1}{mT^{m}}E_{a,b}\left(\frac{1}{T}\right),R_{a,b}(T)+\frac{1}{mT^{m}}E_{a,b}(T)\right)$ (15)
$+P_{c}\left(R_{a,b}(T)+\frac{1}{mT^{m}}E_{a,b}(T),R_{a,b}\left(\frac{1}{T}\right)-\frac{1}{mT^{m}}E_{a,b}\left(\frac{1}{T}\right)\right).$ (16)
@@ 4.2.2 $\max(a,b)=b$
4.3 Case 3: $a\geq 2$ even and $b\geq 1$ odd
While explicitly describing $G_{a,b,m}(T)$ by putting $X=R_{a,b,m}(T),Y=R_{-a,-b,m}(T)$ at
$F_{a,b}(X,Y)=-P_{a}(-Y,-X)-P_{b}(-Y,-X)-P_{c}(-X,-Y)$, we only have to take extra care of $-P_{c}(-X,-Y)$, with others same as in Case 2. Note that $G_{a,b}(T)=T^{-a^{2}}+T^{-b^{2}}-T^{-c^{2}}$. So,
@@ 4.3.2 $\max(a,b)=b$
Department of Mathematical Sciences, Seoul National University, Gwanak-ro 1, Gwankak-gu, Seoul, South Korea 08826
E-mail address: snu1719626@snu.ac.kr
Department of Mathematical Sciences, Seoul National University, Gwanak-ro 1, Gwankak-gu, Seoul, South Korea 08826
E-mail address: dohyeongkim@snu.ac.kr

### v3

@@ 1. Introduction
This is still an open problem. Ennola conjectured in [Enn91] that $\{\epsilon_{l},\epsilon_{l}-1\}$ above is a pair of fundamental units for the maximal order $\mathcal{O}$ of $\mathbb{Q}(\epsilon_{l})$. We call this Ennola’s conjecture. He showed that his conjecture is true for $3\leq l\leq 500$ and that the unit index
@@
To show that Ennola’s conjecture holds true is equivalent to prove $j_{l}=1$ for $\forall l\geq 3$. Louboutin obtained several results on this conjecture. For example, in [Lou17], he showed that for $l\geq 3$, the unit index $j_{l}$ is coprime to $19!$, and $j_{l}=1$ for $3\leq l\leq 5\cdot 10^{7}$. He also proved in [Lou21] that if we assume ABC conjecture is true, then Ennola’s conjecture is true except for finitely many $l$. i.e., $j_{l}=1$ for any sufficiently large $l$. Another significant result in [Lou21] is a conditional proof of a weak form of Ennola’s conjecture that for any given integer $N\geq 2$, we have $\gcd(j_{l},N!)=1$ for $l\geq l_{N}$ effectively large enough.
@@
Theorem 1 is a statement about divisors of a unit index $j_{l}$. A similar work was done by Louboutin and Lee for another family of cubic number fields. In [LL], they showed that the unit index $(\mathbb{U}_{a}:\langle-1,\epsilon_{a},\epsilon_{a}^{\prime}\rangle)$ of the group of units generated by $-1$, $\epsilon_{a}$, and $\epsilon_{a}^{\prime}$ in the group of units $\mathbb{U}_{a}$ of the ring of integers of $\mathbb{Q}(\alpha)$ is coprime to 3 for $a\geq 1$ and coprime to 6 for $1\leq a\equiv 2,3\ (\text{mod }4)$, where $\alpha$ is a root of $X^{3}-4a^{2}X+2\in\mathbb{Z}[X]$, and $\epsilon_{a},\epsilon_{a}^{\prime}$ are two distinct units in the ring of integers of $\mathbb{Q}(\alpha)$. Our result is similar to that of Louboutin and Lee, and we are concerned with a family of cubic polynomials (1) with $l\geq 3$. On the other hand, this family (1) attracts interests both of number theorists and of topologists. For topological context, see [KY].
@@ 2. Louboutin’s conjectures
One can directly compute $P_{d}(X,Y)$ for small $d\geq 1$ by finding all possible pairs of $(k,l)$, and below are some examples of $P_{d}(X,Y)$ obtained from Definition 1.
@@
Then for $d\in\left\{a,b,c\right\}$ and $P_{d}(X,Y)$ as in the Definition 1, we have
(2) $P_{|d|}(S_{a,b}(T),S_{a,b}(1/T))=-S_{a,b}(1/T^{|d|}).$
@@
(3) $P_{|d|}(-R_{a,b}(T),-R_{a,b}(1/T))=\begin{cases}-S_{a,b}(1/T^{|d|}),&\text{if}\ d=a\\ R_{a,b}(1/T^{|d|}),&\text{if}\ d\in\left\{b,c\right\}.\end{cases}$
@@ 3. Proof of Conjecture 1
Hence, we may regard $s_{d}(x_{1},x_{2},x_{3})$ as a polynomial of $\sigma_{1},\sigma_{2},\sigma_{3}$ for each $d\geq 4$. i.e.,
(4) $s_{d}(x_{1},x_{2},x_{3})=x_{1}^{d}+x_{2}^{d}+x_{3}^{d}=f_{d}(\sigma_{1},\sigma_{2},\sigma_{3})$
@@
(5) $f_{d}(\sigma_{1},\sigma_{2},\sigma_{3})=\sigma_{1}f_{d-1}(\sigma_{1},\sigma_{2},\sigma_{3})-\sigma_{2}f_{d-2}(\sigma_{1},\sigma_{2},\sigma_{3})+\sigma_{3}f_{d-3}(\sigma_{1},\sigma_{2},\sigma_{3}).$
@@
(6) $f_{d}(\sigma_{1},\sigma_{2},1)=\sigma_{1}f_{d-1}(\sigma_{1},\sigma_{2},1)-\sigma_{2}f_{d-2}(\sigma_{1},\sigma_{2},1)+f_{d-3}(\sigma_{1},\sigma_{2},1).$
@@
(9) $-f_{d-i}(Y,X,1)=(d-i)\sum_{\begin{subarray}{c}k,l\geq 0\\ 0\leq 2k+3l\leq d-i\end{subarray}}(-1)^{k-1}\binom{k+l}{k}\binom{d-i-k-2l}{k+l}\frac{X^{k}Y^{d-i-2k-3l}}{d-i-k-2l}$
@@
(11) $-\frac{d(d-1-2l)!}{l!\ (d-3l)!}Y^{d-3l}.$
Since we assumed $k=0$, the second term in the case (i), which was obtained by replacing $k$ with $k-1$, cannot appear. Hence,
@@
(12) $(-1)^{k-1}\frac{d(d-1-k)!}{k!\ (d-1-2k)!}X^{k}Y^{d-2k}$
as in the case (ii) by putting $l=0$, and excluding the third term in the case (i). That is,
@@
and the $f_{d}$ comes from Newton identities. Combining (4) with the discussion right after Example 2, we get the desired result. ∎
@@ 4. Proof of Conjecture 2
Table 3. $G_{a,b}(T)$ in each case.
4.1. Case 1: $a\geq 1$ odd and $b\geq 1$ odd
@@
(13) $G_{a,b,m}(T)=F_{a,b}\left(R_{a,b}(T)+\frac{1}{mT^{m}}E_{a,b}(T),R_{-a,-b}(T)+\frac{1}{mT^{m}}E_{-a,-b}(T)\right).$
@@
(14) $-P_{a}\left(R_{a,b}\left(\frac{1}{T}\right)-\frac{1}{mT^{m}}E_{a,b}\left(\frac{1}{T}\right),R_{a,b}(T)+\frac{1}{mT^{m}}E_{a,b}(T)\right)$
(15) $-P_{b}\left(R_{a,b}\left(\frac{1}{T}\right)-\frac{1}{mT^{m}}E_{a,b}\left(\frac{1}{T}\right),R_{a,b}(T)+\frac{1}{mT^{m}}E_{a,b}(T)\right)$
(16) $+P_{c}\left(R_{a,b}(T)+\frac{1}{mT^{m}}E_{a,b}(T),R_{a,b}\left(\frac{1}{T}\right)-\frac{1}{mT^{m}}E_{a,b}\left(\frac{1}{T}\right)\right).$
@@ 4.2.2. $\max(a,b)=b$
4.3. Case 3: $a\geq 2$ even and $b\geq 1$ odd
While explicitly describing $G_{a,b,m}(T)$ by putting $X=R_{a,b,m}(T)$ and $Y=R_{-a,-b,m}(T)$ at $F_{a,b}(X,Y)=-P_{a}(-Y,-X)-P_{b}(-Y,-X)-P_{c}(-X,-Y)$, we only have to take extra care of $-P_{c}(-X,-Y)$, with others same as in Case 2. Note that $G_{a,b}(T)=T^{-a^{2}}+T^{-b^{2}}-T^{-c^{2}}$. So,